# Complex exponential integral means spectra of univalent functions and the Brennan conjecture

### v10

@@ 1. Introduction
More known results on the universal integral means spectrum and related topics can be found in the monograph [26, Chapter VIII] and recent survey [31] given by Hedenmalm and Sola.
@@
Here and later, we say a function $f\in\mathcal{S}$ is an extremal function for $B_{b}(\tau)$ if $\beta_{f}(\tau)=B_{b}(\tau)$. In the recent work [38], we have studied Problem 1.11 for real $\tau$. It has been pointed out that the functional $I:f\rightarrow\beta_{f}(\tau),f\in\mathcal{S}_{q}$, is not continuous on $\mathcal{S}_{q}$ under the locally uniformly convergence topology, see [38, Remark 5.8]. In [38], we introduced and studied the integral means spectrum (IMS) functionals on the Teichmüller spaces. It has been proved in [38] that all IMS functionals on the universal Teichmüller space, the closure of the universal Teichmüller curve, and the universal asymptotic Teichmüller space are continuous. It was also shown in [38] that the real exponential integral means spectrum of any univalent function admitting a quasiconformal extension to $\widehat{\mathbb{C}}$ is strictly less than the universal integral means spectrum. This result means that, for $\tau\in\mathbb{R}$ with $\tau\neq 0$, the extremal functions for $B_{b}(\tau)$ cannot be from the class $\mathcal{S}_{q}$. The first main purpose of this paper is to extend these results to the complex setting. We shall prove that
@@ 4.2. Extremal theory of quasiconformal mappings
It is known that $H_{0}([\mu]_{AT(\Omega)})=h_{0}([\mu]_{T(\Omega)})$. We say some $\widetilde{\mu}\in[\mu]_{AT({\Omega})}$ is non-Strebel extremal in $\mu]_{AT({\Omega})}$ if it satisfies that $H_{0}([\mu]_{AT(\Omega)})=h_{0}([\widetilde{\mu}]_{T(\Omega)})=k_{0}([\widetilde{\mu}]_{T(\Omega)})=\|\widetilde{\mu}\|_{\infty}$. It is easy to see that $H_{0}([\mu]_{AT(\Omega)})=h_{0}([\widetilde{\mu}]_{T(\Omega)})=k_{0}([\widetilde{\mu}]_{T(\Omega)})=h^{*}(\widetilde{\mu})=\|\widetilde{\mu}\|_{\infty}$, if $\widetilde{\mu}$ is non-Strebel extremal in $\mu]_{AT({\Omega})}$. From [56, Lemma 4.1], we know that
@@ 6.1. Certain subclasses of rational functions and the main result of this section
Take $R_{3}(z)=\frac{1}{(1-z)^{2}},$ then $R_{3}^{\prime}(z)=\frac{2}{(1-z)^{3}}$ so that $R_{3}$ has no critical points on $\mathbb{T}$.
@@
Since $f^{\prime}$ has no zeros for $r>r_{f}$, the quantity $|f^{\prime}(re^{i\theta})|^{\tau}$ is well-defined for all $r$ sufficiently close to $1$, and thus the $\limsup$ is meaningful. For each $R\in\mathcal{R}$, since $R$ has only finitely many critical points, we know that there is a constant $r_{R}\in(0,1)$ such that $R^{\prime}(z)\neq 0$ for all $z\in\{z:r_{R}<|z|<1\}$. Hence we can define the generalized integral means spectrum ${\widetilde{\beta}}_{R}$ of $R$ as in (6.8).
@@
Here ${\mathbb{A}}=\max\limits_{1\leq j\leq n}\{|A_{j}|\}$. In (6.2), we have used the following inequality [30, Exercises 1.1.4, Page 11],
@@ 7.1. The first part of remarks
We let $\mathcal{S}_{Q}$ be the class of all univalent functions $f$ which belong to $\mathcal{S}$ and admit a quasiconformal extension to $\widehat{\mathbb{C}}$. It is easy to see that $\mathcal{S}_{q}$ is a proper subset of $\mathcal{S}_{Q}$. For each $\theta\in(-\pi,\pi]$, let $\mathcal{S}_{\theta}$ be the subclass of $\mathcal{S}_{Q}$ which consists of all the functions $f$ satisfying that $\lim\limits_{\Delta\ni z\rightarrow e^{i\theta}}f(z)=\infty$. For each $\theta\in(-\pi,\pi]$, $\mathcal{S}_{\theta}$ is a copy of the universal Teichmüller space. By using [57, Lemma 1] and repeating the arguments in the proof of Theorem 4, we can similarly prove that
@@ 7.3. The third part of remarks
Let $R\in\mathcal{U}_{P}$. If $R$ has at least one critical point on $\mathbb{T}$, then, from Theorem 4, Proposition 7.6, Theorem 1.3, and the fact that $\beta_{R}(\tau)=|\tau|-1$ for $\tau\leq-2$, we know that $R$ does not belong to $\mathcal{S}_{Q}$ (defined as in Remark 7.5), and the Schwarzian derivative $S_{R}$ of $R$ may lie on $\partial\mathbf{S}_{Q}$ (here $\mathbf{S}_{Q}$ is defined as in (7.1)). For example, the function $\mathcal{E}_{2}(z)=z-\frac{1}{2}z^{2}$ has a critical point on $\mathbb{T}$, and we know from [38, Remark 5.9] that $S_{\mathcal{E}_{2}}$ lies on $\partial\mathbf{S}_{Q}$. Hence we believe that $S_{R}$ should lie on the boundary of $\mathbf{S}_{Q}$ if $R\in\mathcal{U}_{R}$ has at least one critical point on $\mathbb{T}$, although we have not found a proof for this claim. We leave it as the following
@@
As remarked above, the proof of part (1) of Conjecture 7.21 is easy. However, checking the validity of part (2) of Conjecture 7.21 seems not easy. We will give an example for part (2) of Conjecture 7.21. We consider again the function $\mathcal{E}_{2}(z)=z-\frac{1}{2}z^{2}$ which has one critical point on $\mathbb{T}$ and belongs to $\partial\mathcal{T}$. It is known from [38, Remark 5.9 $({\bf II})$] that
@@
When $P_{n}\in\mathcal{U}_{P}$ has no critical points on $\mathbb{T}$, we have pointed out that $N_{P_{n}}$ belongs to $E_{1,0}$}. On the other hand, for a univalent function (conformal mapping) $f$ from $\Delta$ to a bounded Jordan domain in $\mathbb{C}$, we know that $f$ is an asymptotically conformal mapping if and only if $N_{f}$ belongs to $E_{1,0}$, see [25], [47] or [49]. Hence it is natural to ask whether $P_{n}$ is an asymptotically conformal mapping in $\Delta$ when $P_{n}\in\mathcal{U}_{P}$ has no critical points on $\mathbb{T}$. However, we find that there exists $P_{n}\in\mathcal{U}_{P}$ with no critical points on $\mathbb{T}$ for which $P_{n}(\mathbb{T})$ is not a bounded Jordan curve so that $P_{n}$ cannot be asymptotically conformal. As an example, we consider ${\bf{P}}(z):=z+\frac{\sqrt{3}}{2}z^{2}+\frac{1}{4}z^{3},z\in\Delta.$ Note that ${\bf{P}}(z)=\pi_{2}\circ\overline{{\bf{P}}}\circ\pi_{1}(z),z\in\Delta$. Here, $w=\pi_{1}(z)=\frac{\sqrt{3}}{2}z$, which maps $\Delta$ conformally to the disk $\Delta(\frac{\sqrt{3}}{2})=\{w\in\mathbb{C}:|w|<\sin(\frac{\pi}{3})=\frac{\sqrt{3}}{2}\}$, $\zeta=\overline{{\bf{P}}}(w)=(1+w)^{3}-1$, $\pi_{2}(\zeta)=\frac{2}{3\sqrt{3}}\zeta$. Then, by Corollary 5.7.3 in [52, Page 197], we obtain that $\bf{P}$} belongs to $\mathcal{U}_{P}$, or we can check this fact by using Dieudonné’s univalence criterion, see [19, Page 75]. Meanwhile, we see from ${\bf{P}}^{\prime}(z)=(\frac{\sqrt{3}}{2}z+1)^{2}$ that ${\bf{P}}^{\prime}\neq 0$ for any $|z|<2/\sqrt{3}$. Also, we easily check that ${\bf{P}}(-\frac{\sqrt{3}}{2}+\frac{1}{2}i)={\bf{P}}(-\frac{\sqrt{3}}{2}-\frac{1}{2}i)$, which means that the curve ${\bf P}(\mathbb{T})$ is not a bounded Jordan curve in $\mathbb{C}$.
@@ 7.4. The final part of remarks
It is easy to see that, if for any $\alpha>0$, (7.4) holds for all $f\in\mathcal{S}_{q}$, then the Brennan conjecture is true. On the other hand, it has been pointed out in [53, Page 1633] that there exists univalent function $f\in\mathcal{S}$ such that $\mathbb{M}_{f}(\alpha)>\mathbb{M}_{\kappa}(\alpha)$ when $\alpha>-1$. Meanwhile, for any $\alpha>-1$, it has been shown in [53] that,
@@
$\frac{\Gamma(n+\lambda)}{n!}=n^{\lambda-1}[1+o(1)],\,\,{\textup{as}}\,\,n\rightarrow\infty.$
@@
(7.10) $\|{(1-rz)^{-\lambda}}\|_{\alpha}^{2}$ $=$ $\sum_{n=0}^{\infty}\frac{\Gamma(\alpha+2)}{n^{\alpha+1}}[1+o(1)]\cdot\frac{n^{2\lambda-2}}{[\Gamma(\lambda)]^{2}}[1+o(1)]r^{2n}$
$=$ $\sum_{n=0}^{\infty}\frac{\Gamma(\alpha+2)}{[\Gamma(\lambda)]^{2}}\cdot{n^{2\lambda-\alpha-3}}[1+o(1)]r^{2n}$
@@
for any $r\in(0,1)$, see [53, Page 1631], hence, we see that, for $2\lambda>\alpha+2$,
@@
We conjecture that
Conjecture 7.33.
For every $\alpha>0$, if the functional $\Lambda_{\alpha}$ attains at least one maximum on $S_{f}$ in $\overline{\mathbf{S}}$ for some $f\in\mathcal{S}$, then $\|S_{f}\|_{E_{2}}=6$.
@@
In light of Proposition 7.26, to find more functions satisfying the Brennan conjecture, it is natural to study the problem that in which subset of $\mathcal{S}_{Q}$ the Koebe function and its rotations provide the biggest norm for the multiplication operator when $\alpha>0$. That is
Problem 7.35.
For each $\alpha>0$, which functions $f\in\mathcal{S}_{Q}$ satisfy $\mathbb{M}_{f}(\alpha)\leq\mathbb{M}_{\kappa}(\alpha)$?
@@
Let $\alpha>0,f\in\mathcal{S}$. By checking carefully the arguments in the proof of Proposition $8$ in [53, Page 1632], we see that, if there is a constant $\varepsilon>0$ such that
@@
We conclude by remarking that, if Problem 7.35 has an affirmative answer for every $f\in\mathcal{S}_{Q}$, i.e., $\mathbb{M}_{f}(\alpha)\leq\mathbb{M}_{\kappa}(\alpha)$ holds for all $\alpha>0$ and all $f\in\mathcal{S}_{Q}$, then the Brennan conjecture follows. Furthermore, since Claim 7.37 holds, then it follows that $\mathbb{M}_{R}(\alpha)=\mathbb{M}_{\kappa}(\alpha)$ for every $\alpha>0$ and all $R\in\mathcal{U}_{R}$ that have at least one critical point on $\mathbb{T}$. Although these predictions may seem surprising, no counterexamples are known to the author.

### v11

@@ 1. Introduction
More known results on the universal integral means spectrum and related topics can be found in the monograph [27, Chapter VIII] and recent survey [32] given by Hedenmalm and Sola.
@@
Here and later, we say a function $f\in\mathcal{S}$ is an extremal function for $B_{b}(\tau)$ if $\beta_{f}(\tau)=B_{b}(\tau)$. In the recent work [39], we have studied Problem 1.11 for real $\tau$. It has been pointed out that the functional $I:f\rightarrow\beta_{f}(\tau),f\in\mathcal{S}_{q}$, is not continuous on $\mathcal{S}_{q}$ under the locally uniformly convergence topology, see [39, Remark 5.8]. In [39], we introduced and studied the integral means spectrum (IMS) functionals on the Teichmüller spaces. It has been proved in [39] that all IMS functionals on the universal Teichmüller space, the closure of the universal Teichmüller curve, and the universal asymptotic Teichmüller space are continuous. It was also shown in [39] that the real exponential integral means spectrum of any univalent function admitting a quasiconformal extension to $\widehat{\mathbb{C}}$ is strictly less than the universal integral means spectrum. This result means that, for $\tau\in\mathbb{R}$ with $\tau\neq 0$, the extremal functions for $B_{b}(\tau)$ cannot be from the class $\mathcal{S}_{q}$. The first main purpose of this paper is to extend these results to the complex setting. We shall prove that
@@ 4.2. Extremal theory of quasiconformal mappings
It is known that $H_{0}([\mu]_{AT(\Omega)})=h_{0}([\mu]_{T(\Omega)})$. We say some $\widetilde{\mu}\in[\mu]_{AT({\Omega})}$ is non-Strebel extremal in $\mu]_{AT({\Omega})}$ if it satisfies that $H_{0}([\mu]_{AT(\Omega)})=h_{0}([\widetilde{\mu}]_{T(\Omega)})=k_{0}([\widetilde{\mu}]_{T(\Omega)})=\|\widetilde{\mu}\|_{\infty}$. It is easy to see that $H_{0}([\mu]_{AT(\Omega)})=h_{0}([\widetilde{\mu}]_{T(\Omega)})=k_{0}([\widetilde{\mu}]_{T(\Omega)})=h^{*}(\widetilde{\mu})=\|\widetilde{\mu}\|_{\infty}$, if $\widetilde{\mu}$ is non-Strebel extremal in $\mu]_{AT({\Omega})}$. From [57, Lemma 4.1], we know that
@@ 6.1. Certain subclasses of rational functions and the main result of this section
Take $R_{3}(z)=\frac{z-\frac{1}{2}z^{2}}{(1-z)^{2}},$ then $R_{3}^{\prime}(z)=\frac{1}{(1-z)^{3}}$ so that $R_{3}$ has no critical points on $\mathbb{T}$.
@@
Since $f^{\prime}(re^{i\theta})$ has no zeros for $r$ satisfying $r_{f}<r<1$, the quantity $|f^{\prime}(re^{i\theta})|^{\tau}$ is well-defined for all $r$ sufficiently close to $1$, and thus the $\limsup$ is meaningful and ${\widetilde{\beta}}_{f}$ can be equivalently defined as
(6.9) ${\widetilde{\beta}}_{f}(\tau)=\lim\limits_{r_{f}<\gamma\to 1^{-}}\sup\limits_{r\in(\gamma,1)}\frac{\log\int_{-\pi}^{\pi}|f^{\prime}(re^{i\theta})|^{\tau}d\theta}{|\log(1-r)|}.$
For each $R\in\mathcal{R}$, since $R$ has only finitely many critical points, we know that there is a constant $r_{R}\in(0,1)$ such that $R^{\prime}(z)\neq 0$ for all $z\in\{z:r_{R}<|z|<1\}$. Hence we can define the generalized integral means spectrum ${\widetilde{\beta}}_{R}$ of $R$ as in (6.8) or (6.9).
@@
Here ${\mathbb{A}}=\max\limits_{1\leq j\leq n}\{|A_{j}|\}$. In (6.2), we have used the following inequality [31, Exercises 1.1.4, Page 11],
@@
Remark 6.27.
We recall the Schwarz-Christoffel formula for the unit disk, see [1, Page 236]. Let $f$ be a univalent function (conformal map) from the unit disk $\Delta$ onto the interior of a bounded polygon $\mathcal{P}$ with vertices $z_{1},\dots,z_{n},n\geq 3$, and corresponding interior angles $\alpha_{1}\pi,\dots,\alpha_{n}\pi$. Then the derivative of $f$ is given by
$f^{\prime}(z)=C\prod_{k=1}^{n}\left(z-{a_{k}}\right)^{\alpha_{k}-1},\quad z\in\Delta,$
where $a_{1},a_{2},\dots,a_{n}$ are points on the unit circle $\mathbb{T}$ and $C$ is a non-zero complex constant.
Note that for $r\in(0,1)$,
$\int_{-\pi}^{\pi}\frac{d\theta}{|f^{\prime}(re^{i\theta})|^{2}}=C^{-2}\int_{-\pi}^{\pi}\frac{d\theta}{\prod\limits_{k=1}^{n}\left|re^{i\theta}-{e^{i\arg a_{k}}}\right|^{2\alpha_{k}-2}}.$
Since $\alpha_{k}<2$ for each $k$, a direct estimate using Lemma 6.22 and similar arguments in the proof of part $(\bf c_{3})$ of Theorem 6.18 shows that $\beta_{f}(-2)<1$. Actually, let
${\bf{\pi}}(z)={\bf{\pi}}(re^{i\theta})=\frac{1}{\prod\limits_{k=1}^{n}\left|re^{i\theta}-{e^{i\arg a_{k}}}\right|^{2\alpha_{k}-2}}.$
For a given $r_{0}\in(0,1)$, we can find constants $M>0$, $\eta_{k}\in(0,\frac{\pi}{2n})$, $k=1,2,\cdots,n$ such that $\Omega_{k_{1}}\cap\Omega_{k_{2}}=\emptyset$ whenever $k_{1}\neq k_{2}$, $1\leq k_{1},k_{2}\leq n,$ here $\Omega_{k}=\{re^{\theta}:r_{0}\leq r\leq 1,\arg a_{k}-\eta_{k}\leq\theta\leq\arg a_{k}+\eta_{k}\},$ and $|{\bf{\pi}}(z)|\leq M$ for all $z\in\Omega-\bigcup\limits_{1\leq k\leq n}\Omega_{k}$, here $\Omega=\{re^{i\theta}:r_{0}\leq r\leq 1\}$.
For $1\leq k\leq n$, we take
$D^{k}:=\max\limits_{z\in\Omega_{k}}\{\prod\limits_{1\leq b\leq n,b\neq k}|z-e^{i\arg a_{k}}|^{2-2\alpha_{k}}\}.$
We define ${\bf\mathfrak{D}}^{\ast}=\max\limits_{1\leq k\leq n}\{D_{k}\}.$ Then we easily see that $0<\mathfrak{D}^{\ast}<+\infty$ since $\Omega_{k},k=1,2,\cdots,n$ are all compact. Hence, when $r>r_{0}$, we have
$\int_{-\pi}^{\pi}\pi(re^{i\theta})d\theta$ $\leq$ $\sum\limits_{k=1}^{n}\int_{\arg a_{k}-\eta_{k}}^{\arg a_{k}+\eta_{k}}\pi(re^{i\theta})d\theta+2\pi M$
$\leq$ $[\mathfrak{D}^{\ast}]^{2}\sum\limits_{j=1}^{n}\int_{\arg a_{k}-\eta_{k}}^{\arg a_{k}+\eta_{k}}\frac{d\theta}{|re^{i\theta}-e^{i\arg a_{k}}|^{2\alpha_{k}-2}}+2\pi M$
$=$ $[\mathfrak{D}^{\ast}]^{2}\sum\limits_{k=1}^{n}\int_{-\eta_{k}}^{\eta_{k}}\frac{d\theta}{|1-re^{i\theta}|^{2\alpha_{k}-2}}+2\pi M.$
Consequently, from Lemma 6.22 and the fact that $\alpha_{k}<2$ for every $k$, we conclude that $\beta_{f}(-2)<1$. That is to say, the Brennan conjecture holds for all univalent functions mapping $\Delta$ onto the interior of a bounded polygon $\mathcal{P}$.
@@ 7.1. The first part of remarks
We let $\mathcal{S}_{Q}$ be the class of all univalent functions $f$ which belong to $\mathcal{S}$ and admit a quasiconformal extension to $\widehat{\mathbb{C}}$. It is easy to see that $\mathcal{S}_{q}$ is a proper subset of $\mathcal{S}_{Q}$. For each $\theta\in(-\pi,\pi]$, let $\mathcal{S}_{\theta}$ be the subclass of $\mathcal{S}_{Q}$ which consists of all the functions $f$ satisfying that $\lim\limits_{\Delta\ni z\rightarrow e^{i\theta}}f(z)=\infty$. For each $\theta\in(-\pi,\pi]$, $\mathcal{S}_{\theta}$ is a copy of the universal Teichmüller space. By using [58, Lemma 1] and repeating the arguments in the proof of Theorem 4, we can similarly prove that
@@ 7.3. The third part of remarks
Let $R\in\mathcal{U}_{P}$. If $R$ has at least one critical point on $\mathbb{T}$, then, from Theorem 4, Proposition 7.6, Theorem 1.3, and the fact that $\beta_{R}(\tau)=|\tau|-1$ for $\tau\leq-2$, we know that $R$ does not belong to $\mathcal{S}_{Q}$ (defined as in Remark 7.5), and the Schwarzian derivative $S_{R}$ of $R$ may lie on $\partial\mathbf{S}_{Q}$ (here $\mathbf{S}_{Q}$ is defined as in (7.1)). For example, the function $\mathcal{E}_{2}(z)=z-\frac{1}{2}z^{2}$ has a critical point on $\mathbb{T}$, and we know from [39, Remark 5.9] that $S_{\mathcal{E}_{2}}$ lies on $\partial\mathbf{S}_{Q}$. Hence we believe that $S_{R}$ should lie on the boundary of $\mathbf{S}_{Q}$ if $R\in\mathcal{U}_{R}$ has at least one critical point on $\mathbb{T}$, although we have not found a proof for this claim. We leave it as the following
@@
As remarked above, the proof of part (1) of Conjecture 7.21 is easy. However, checking the validity of part (2) of Conjecture 7.21 seems not easy. We will give an example for part (2) of Conjecture 7.21. We consider again the function $\mathcal{E}_{2}(z)=z-\frac{1}{2}z^{2}$ which has one critical point on $\mathbb{T}$ and belongs to $\partial\mathcal{T}$. It is known from [39, Remark 5.9 $({\bf II})$] that
@@
When $P_{n}\in\mathcal{U}_{P}$ has no critical points on $\mathbb{T}$, we have pointed out that $N_{P_{n}}$ belongs to $E_{1,0}$}. On the other hand, for a univalent function (conformal mapping) $f$ from $\Delta$ to a bounded Jordan domain in $\mathbb{C}$, we know that $f$ is an asymptotically conformal mapping if and only if $N_{f}$ belongs to $E_{1,0}$, see [26], [48] or [50]. Hence it is natural to ask whether $P_{n}$ is an asymptotically conformal mapping in $\Delta$ when $P_{n}\in\mathcal{U}_{P}$ has no critical points on $\mathbb{T}$. However, we find that there exists $P_{n}\in\mathcal{U}_{P}$ with no critical points on $\mathbb{T}$ for which $P_{n}(\mathbb{T})$ is not a bounded Jordan curve so that $P_{n}$ cannot be asymptotically conformal. As an example, we consider ${\bf{P}}(z):=z+\frac{\sqrt{3}}{2}z^{2}+\frac{1}{4}z^{3},z\in\Delta.$ Note that ${\bf{P}}(z)=\pi_{2}\circ\overline{{\bf{P}}}\circ\pi_{1}(z),z\in\Delta$. Here, $w=\pi_{1}(z)=\frac{\sqrt{3}}{2}z$, which maps $\Delta$ conformally to the disk $\Delta(\frac{\sqrt{3}}{2})=\{w\in\mathbb{C}:|w|<\sin(\frac{\pi}{3})=\frac{\sqrt{3}}{2}\}$, $\zeta=\overline{{\bf{P}}}(w)=(1+w)^{3}-1$, $\pi_{2}(\zeta)=\frac{2}{3\sqrt{3}}\zeta$. Then, by Corollary 5.7.3 in [53, Page 197], we obtain that $\bf{P}$} belongs to $\mathcal{U}_{P}$, or we can check this fact by using Dieudonné’s univalence criterion, see [20, Page 75]. Meanwhile, we see from ${\bf{P}}^{\prime}(z)=(\frac{\sqrt{3}}{2}z+1)^{2}$ that ${\bf{P}}^{\prime}\neq 0$ for any $|z|<2/\sqrt{3}$. Also, we easily check that ${\bf{P}}(-\frac{\sqrt{3}}{2}+\frac{1}{2}i)={\bf{P}}(-\frac{\sqrt{3}}{2}-\frac{1}{2}i)$, which means that the curve ${\bf P}(\mathbb{T})$ is not a bounded Jordan curve in $\mathbb{C}$.
@@ 7.4. The final part of remarks
It is easy to see that, if for any $\alpha>0$, (7.4) holds for all $f\in\mathcal{S}_{q}$, then the Brennan conjecture is true. On the other hand, it has been pointed out in [54, Page 1633] that there exists univalent function $f\in\mathcal{S}$ such that $\mathbb{M}_{f}(\alpha)>\mathbb{M}_{\kappa}(\alpha)$ when $\alpha>-1$. Meanwhile, for any $\alpha>-1$, it has been shown in [54] that,
@@
$\frac{\Gamma(n+\lambda)}{n!}=(n+1)^{\lambda-1}[1+o(1)],\,\,{\textup{as}}\,\,n\rightarrow\infty.$
@@
(7.10) $\|{(1-rz)^{-\lambda}}\|_{\alpha}^{2}$ $=$ $\sum_{n=0}^{\infty}\frac{\Gamma(\alpha+2)}{(n+1)^{\alpha+1}}[1+o(1)]\cdot\frac{(n+1)^{2\lambda-2}}{[\Gamma(\lambda)]^{2}}[1+o(1)]r^{2n}$
$=$ $\sum_{n=0}^{\infty}\frac{\Gamma(\alpha+2)}{[\Gamma(\lambda)]^{2}}\cdot{(n+1)^{2\lambda-\alpha-3}}[1+o(1)]r^{2n}$
@@
for any $r\in(0,1)$, see [54, Page 1631], hence, we see that, for $2\lambda>\alpha+2$,
@@
In light of Proposition 7.26, to find more functions satisfying the Brennan conjecture, it is natural to study the problem that in which subset of $\mathcal{S}$ the Koebe function and its rotations provide the biggest norm for the multiplication operator when $\alpha>0$. That is
Problem 7.34.
For each $\alpha>0$, which functions $f\in\mathcal{S}$ satisfy $\mathbb{M}_{f}(\alpha)\leq\mathbb{M}_{\kappa}(\alpha)$?
@@
Let $\alpha>0,f\in\mathcal{S}$. By checking carefully the arguments in the proof of Proposition $8$ in [54, Page 1632], we see that, if there is a constant $\varepsilon>0$ such that
@@
If our guess that Problem 7.34 has an affirmative answer for every $f$ in $\mathcal{S}$ is true, i.e., $\mathbb{M}_{f}(\alpha)\leq\mathbb{M}_{\kappa}(\alpha)$ holds for every $\alpha>0$ and all $f\in\mathcal{S}$, then the Brennan conjecture follows. Furthermore, since Claim 7.36 holds, it follows that $\mathbb{M}_{R}(\alpha)=\mathbb{M}_{\kappa}(\alpha)$ for every $\alpha>0$ and all $R\in\mathcal{U}_{R}$ that have at least one critical point on $\mathbb{T}$. Although these predictions may seem surprising, no counterexamples are known to this author. Finally, based on the above findings, we further conjecture that
Conjecture 7.40.
For every $\alpha>0$, if the functional $\Lambda_{\alpha}$ attains the maximum on $S_{f}$ in $\overline{\mathbf{S}}$ for some $f\in\mathcal{S}$, then $\|S_{f}\|_{E_{2}}=6$.